what is the answer to this problem: ((-2)/(c^2+2c-3))-((5)/(3-3c))=((4)/(3c+9))
im really having a lot of trouble with it.

Answer by lwsshak3(11628)

what is the answer to this problem: ((-2)/(c^2+2c-3))-((5)/(3-3c))=((4)/(3c+9))
im really having a lot of trouble with it.
..
((-2)/(c^2+2c-3))-((5)/(3-3c))=((4)/(3c+9))
-2/(c+3)(c-1)-5/3(1-c)=4/3(c+3)
2/(c+3)(1-c)-5/3(1-c)=4/3(c+3)
LCD:3(c+3)(1-c)
2*3-5(c+3)=4(1-c)
6-5c-15=4-4c
c=-13
Check:
((-2)/(c^2+2c-3))-((5)/(3-3c))=((4)/(3c+9))
-2/(169-26-3)-5/(3+39)=4/(-39+9)
-.0143-.1190=-.1333